OpenStudy (anonymous):

Please help!!! (algebra) "Given the system of equations, what is the value of the system determinant? 2x + y = 8 x - y = 10"

OpenStudy (anonymous):

Okay, so I know that to find the system determinant I need to make a matrix...but I'm confused. Don't I need at least 3 variables to do that? What exactly is a "system determinant"...how could I use it in a problem? Most importantly, how do I solve for the system determinant?

OpenStudy (campbell_st):

ok... so if you add your 2 equations y is eliminated... and you will be able to solve for x. when you have x, substitute it into either equation... and find y

OpenStudy (anonymous):

allright x = 4 y = -6

OpenStudy (campbell_st):

not quite add the equations 2x + y = 8 x - y = 10 ---------- 3x = 18 solve for x...

OpenStudy (anonymous):

ok, x = 6 y = -4

OpenStudy (campbell_st):

that looks a bit better

OpenStudy (anonymous):

is it correct?

OpenStudy (campbell_st):

I'd say yes...

OpenStudy (campbell_st):

you can check by substituting beck into each equation.

OpenStudy (anonymous):

ok, now how do i find the system determinant?

OpenStudy (anonymous):

you have a perfect guy here. let him help

OpenStudy (anonymous):

@campbell_st "using determinants"

OpenStudy (anonymous):

when you have equations\[a_1x+b_1y=c_1\qquad\text{and}\\a_2x+b_2y=c_2\] then, the system can be represented in a matrix form as \[\left[\begin{matrix} a_1&b_1\\a_2&b_2 \end{matrix}\right]\left[\begin{matrix} x\\y \end{matrix}\right]=\left[\begin{matrix} c_1\\c_2 \end{matrix}\right] \]

OpenStudy (anonymous):

now, write your system of equations in this format

OpenStudy (anonymous):

@ScienceQs

OpenStudy (anonymous):

I only have x and y...how do I get "b" and "a"?

OpenStudy (anonymous):

\[(2)x+(1)y=8\\(1)x+(-1)y=10\]

OpenStudy (anonymous):

get it?

OpenStudy (anonymous):

ok, but why do I do that? will I multiply 2 times x like that every time I solve a problem tp find the system determinant?

OpenStudy (anonymous):

nobody said to multiply.. just write them down

OpenStudy (anonymous):

this matrix determinant is the easiest and ONLY method to solve large problems.

OpenStudy (anonymous):

smaller ones you can getaway with the method shown by campbell above

OpenStudy (anonymous):

now, lets get back to creating our matrices

OpenStudy (anonymous):

ok, so i wrote it down. now what? oh, and should I substitute x and y for their real values in the matrix?

OpenStudy (anonymous):

nope.. hold your horses

OpenStudy (anonymous):

so, it would look like

OpenStudy (anonymous):

\[\left[\begin{matrix} 2&1\\1&-1 \end{matrix}\right]\left[\begin{matrix} x\\y \end{matrix}\right]=\left[\begin{matrix} 8\\10 \end{matrix}\right] \]

OpenStudy (anonymous):

now, think of matrices like numbers \[AX=B\] so, \[{1\over A}AX={1\over A}B\]

OpenStudy (anonymous):

we cool?

OpenStudy (anonymous):

@ScienceQs ola

OpenStudy (anonymous):

I see how you change a linear equation into a matrix now, but I am still confused... does A = 2 and B = 1, always? or does A and B represent a matrix? If so, which one?

OpenStudy (anonymous):

no. A represents the entire first block with numbers (2 by 2) X represents the second block (with variables in it -> we need to find this) C represents the last block (again of numbers but smaller)

OpenStudy (anonymous):

i meant to say B, not C

OpenStudy (anonymous):

k, so A =[211−1]?

OpenStudy (anonymous):

bingo

OpenStudy (anonymous):

now look at that "A" dividin.. how can we divide with a "block" of numbers?

OpenStudy (anonymous):

yay! and X = [xy]?

OpenStudy (anonymous):

yep. but the format is important... coming back to our topic. did you understand my recent comments?

OpenStudy (anonymous):

so, we are wanting to divide A? How do we do that?

OpenStudy (anonymous):

yes. getting there

OpenStudy (anonymous):

instead of dividing, we convert it to multiplication... like this: \[{1\over A}=A^{-1}\] the \(A^{-1}\) is called the inverse of the matrix \(A\)

OpenStudy (anonymous):

so, what we now have is: \[A^{-1}AX=A^{-1}B\] but, \(A^{-1}A=I\) I is the identity matrix.. it has similar properties like our regular number "1" but using a bulk (matrix)

OpenStudy (anonymous):

\[[12 -11]\] = A^-1 then?

OpenStudy (anonymous):

so, what we now have is \[X=A^{-1}B\] yay!

OpenStudy (anonymous):

nope.. again, keep holding your horses.. dont let em loose yet

OpenStudy (anonymous):

understand that since A is a group of numbers, so is its inverse.

OpenStudy (anonymous):

ok! i don't really understand much of that, but let's move on. what's next?

OpenStudy (anonymous):

if\[A=\left[\begin{matrix} a_1&b_1\\a_2&b_2 \end{matrix}\right]\] then \[A^{-1}=\frac{1}{a_1b_2-a_2b_1} \left[\begin{matrix} b_2&-b_1\\-a_2&b_1 \end{matrix}\right]\]

OpenStudy (anonymous):

we good? keep the horses ready

OpenStudy (anonymous):

ok, i think I got that about A

OpenStudy (anonymous):

now, plug in the numbers.. \[ A^{-1}=\frac{1}{(2)(-1)-(1)(1)} \left[\begin{matrix} -1&-1\\-1&1 \end{matrix}\right]\\ \therefore A^{-1}=\frac{1}{-3} \left[\begin{matrix} -1&-1\\-1&1 \end{matrix}\right]\\\therefore X=\frac{1}{-3} \left[\begin{matrix} -1&-1\\-1&1 \end{matrix}\right] \left[\begin{matrix} 8\\10 \end{matrix}\right] \]

OpenStudy (anonymous):

and now you let em loose

OpenStudy (anonymous):

the three dots, those mean "therefore" right? and the -3 and the -1s...how do I know when to put those in? Like, would it work the same if I put in a -2 where there is a -1? How do I know when to use -1...I am very confused/curious about that.

OpenStudy (anonymous):

yes..(therefore) each line is a step... did you see how I put the numbers for find the inverse?

OpenStudy (anonymous):

compare the original A with the formula for inverse

OpenStudy (anonymous):

you do not get to change the signs... just plug the numbers with their signs and the numbers will reveal themselves

OpenStudy (anonymous):

do you think you could refer me to a good webpage about matrices and how they work? i have a lesson on it but maybe there's another way to explain it that I would get a lot better

OpenStudy (anonymous):

all hail wikipeda

OpenStudy (anonymous):

sincerely, only a one-on-one tutor can teach you any better than this!

OpenStudy (anonymous):